sir, can you please give me all the methods used for finding equivalent resistance?

sir,


can you please give me all the methods used for finding equivalent resistance?

Grade:12

2 Answers

Askiitians_Expert Yagyadutt
13 years ago

Hello shashank

Brother basiclly you required only two method to find the equivalent resistance

1) When resistance are parallel ...suppose  R1 and R2

The equivalent resistance is given by

1/R = 1/R1 + 1/R2

2) When resistance are series ..

Then equivalent resistance is given by

R = R1 + R2

Now the fact is you need to find out in the network which resistances are in parallel and which are in series ...so it requires just a bit more practice...Solve atleast 50 problem based on this quivalent resistance problem...then you will get a better method of your own how to find..

Any complex netwoork can be reduce in series and parallel combination of resistance ..and then they can be individually solved

Take my suggestion....make a collection of all the past year problems of IIT and AIEEE.....sove them ....

As much as u will practice in this case..you will get more and more better tricks to solve

 

With regards

Yagya

askiitians_expert

mohit
6 years ago
Hey finding equivalent resistance is quite easy. Only you have to take care about the connecting wires and the junctions in which branches are attached.